OpenStudy (anonymous):

Can someone help explain this? In a certain clothing store, 3 shirts and 5 ties cost $60, and 2 shirts and 3 ties cost $39. What is the cost of each shirt?

OpenStudy (anonymous):

Gorgeous picture from Google Images.

OpenStudy (anonymous):

let the value of shirt=x rs. and the value of tie=y so equation become 3x+5y=60....................(1) 2x+3y=39....................(2) solve this equation and get the ans

OpenStudy (anonymous):

thanks a lot @rizzu :D